OpenStudy (anonymous):

solve each system by elimination 6x+8y=-16 7x+16y=-12 and show work please

OpenStudy (kropot72):

If you multiply both sides of the first equation by 2 it will be possible to eliminate the y terms by subtraction. 12x + 16y = -32 7x + 16y = -12 Subtracting the second equation from the first we get 5x = -20 Now you can solve for x.

OpenStudy (anonymous):

8y=-6x-16 y=-6/8x-6 or 3/4x - 2 7x+16(3/4x - 2)=12 7x+12x-32=12 19x=44 x=44/19

OpenStudy (anonymous):

how would you solve for y

OpenStudy (anonymous):

And put numbers and explain it good because this part confuses me

OpenStudy (kropot72):

The value of x is -4. We can chose either of the original equations an plug in -4 in place of x. Lets take the first equation, 6x+8y=-16 Putting -4 in place of x we get \[(6\times -4)+8y=-16\] which simplifies to \[-24+8y=-16..............(1)\] Now can you add 24 to both sides of equation (1) to eliminate the -24 term from the left hand side?

OpenStudy (anonymous):

Can you explain to me in numbers please

OpenStudy (kropot72):

What does the following equation simplify to? -24 + 24 + 8y = -16 + 24

OpenStudy (anonymous):

I mean explain how to find y in numbers

OpenStudy (kropot72):

I have just given the first stage of finding the value of y. Can you simplify the following equation which has one term in y and the remaining terms are numbers? -24 + 24 + 8y = -16 + 24

OpenStudy (anonymous):

can you go through the steps of the processs to find y totally because I am confused

OpenStudy (kropot72):

Sure. What is the result of the following: -24 + 24 = ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

am I correct

OpenStudy (kropot72):

-24 + 24 = 0 Now can you try the following: -16 + 24 = ?